# Workflow as a Service Broker in Cloud Environment: A Systematic Literature Review

### v1

@@ 4.4. Monitoring and fault tolerance taxonomy
The final method is replication, which involves running duplicate instances of each task across multiple VMs to enhance the reliability of task execution (Taghinezhad-Niar and Taheri, 2023). In the event of a failure, the task can continue running without disruption, as its duplicate instances remain operational on other resources. While this approach ensures task continuity without losing time, it adds additional costs for the broker due to the need for extra resources to run the duplicated tasks. Chen et al. (Chen et al., 2024b) combine replication and re-execution for workflows running on a mix of on-demand and spot instances. They schedule tasks on spot instances only if there is sufficient time for re-execution in case the resource is reclaimed by the provider. Otherwise, the task is scheduled on an on-demand instance. When tasks are assigned to spot instances, they attempt to identify idle time on on-demand instances to replicate the tasks where possible. If a spot instance is reclaimed, tasks running on it are checked for duplicates. Tasks with duplicates are ignored, while those without duplicates are re-executed on another instance.
5. Future directions
Although WaaS brokers and the MWSP have made significant progress, there is still considerable potential for improving scheduling and provisioning algorithms. Specifically, there are areas that have received less attention and could benefit from further development, as well as new technologies, such as serverless computing, that can be leveraged. In the following, we outline some promising approaches for enhancing WaaS brokers, organized by the key components of a WaaS broker.
5.1. Negotiation and Pricing
Negotiation and pricing have not yet been sufficiently explored and remain open research challenges. Researchers could develop new negotiation algorithms to facilitate the creation of SLAs between brokers and users.The only prior work on negotiation mechanisms (Taghavi et al., 2023) involves receiving the user’s workflow deadline and proposing a price, which the user can accept or reject. Similar approaches can be explored for other constraints. For instance, in a budget-constrained scenario, the negotiation component could receive the user’s maximum budget and provide an estimated execution time for the workflow, which could then form the basis for an SLA. In multi-criteria optimization scenarios, the broker might identify several Pareto-optimal solutions and offer them to the user for selection. For example, in balancing execution time and cost minimization, the broker could present a list of non-inferior $\langle time,cost\rangle$ pairs for the user to choose from. The broker could also offer various options for QoS criteria, each associated with different prices, allowing users to select higher-cost, higher-reliability options for more critical or sensitive workflows. Additionally, dynamic negotiation protocols could be explored to create more user-friendly agreements.
Pricing policies also remain underexplored, presenting significant opportunities for further research. Particularly when the broker utilizes reserved or spot instances, pricing policies should incentivize users to submit workflows during off-peak times, when reserved instances are idle and spot instances are more affordable. Policies could also encourage users to set more flexible deadlines, allowing the broker’s scheduler to make better use of these cheaper resources. Furthermore, existing work in this area primarily employs cost-based pricing, where the total workflow execution cost is used to determine the price. Future research could explore more advanced pricing strategies, such as value-based or market-based pricing models (Wu et al., 2019), which are prevalent in economics and cloud computing, as potential foundations for more innovative pricing policies.
5.2. Task Scheduling
In the case of scheduling algorithms, while heuristic and meta-heuristic approaches have reached a certain level of maturity, other methods still hold significant potential for creating more effective scheduling algorithms. Game theory has been employed in only one previous study, yet it seems to have untapped potential for both scheduling and pricing algorithms. Additionally, mechanism design, a branch of game theory focused on designing rules that incentivize agents to act in ways that align with the system’s global objectives, has been successfully applied in scheduling multiple functions (Moakhar and Abrishami, 2024) and workflows (Mahdizadeh and Abrishami, 2024) in edge computing environments. This method could be a promising approach for tackling the MWSP. Furthermore, RL has recently gained significant attention in the scientific community. Although seven previous studies have already applied RL to solve the MWSP, it still offers great potential to propose better solutions to this problem.
Regarding task concurrency, only eight studies have explored running multiple tasks on a single resource. For a WaaS broker tasked with executing tens or even hundreds of tasks concurrently, it is impractical to rent a separate resource for each task. Therefore, the use of container technology to run multiple tasks on a single VM warrants more attention from researchers. Finally, for scheduling types, combining static planning with dynamic scheduling appears to be a very promising approach, as it leverages the strengths of both scheduling strategies. However, only two previous works (Taheri et al., 2023; Zhou et al., 2017) have employed this combination. Other researchers could explore this approach to develop more effective solutions.
5.3. Resource Provisioning
Since the WaaS broker aims to maximize its profits, it is logical to adopt various resource pricing models to minimize its costs. One effective approach is the combination of inexpensive but unreliable spot instances with more costly yet reliable on-demand instances, a strategy employed by six previous studies. However, there is still room for the development of new brokers that leverage this combination. Another promising option is the use of cheaper but non-elastic reserved instances in conjunction with elastic but more expensive on-demand resources. Surprisingly, only two prior works have explored this promising combination, and it is expected that more researchers will consider such models in the future. Additionally, combining all three pricing models—spot, reserved, and on-demand—has the potential to reduce costs while maintaining acceptable levels of reliability, yet none of the existing works have considered this comprehensive combination.
@@ 5.3. Resource Provisioning
The emergence of serverless computing and the FaaS model offers new opportunities for WaaS brokers (Wen et al., 2023). This model eliminates the need for users and brokers to manage resources and handle scheduling. Instead, the cloud provider takes care of provisioning, scaling, and function scheduling. Additionally, FaaS operates on a utilization-based, fine-grained pricing system, charging users only for the actual resources they consume, which resolves concerns about idle resources. Although the application of FaaS for scientific workflows has been explored by some researchers (Malawski et al., 2020), the use of serverless platforms in workflow execution is still in its early stages. Most existing research focuses on frameworks designed to run a single workflow in a serverless environment. Notable examples include StepConf (Wen et al., 2022), Wukong (Carver et al., 2020), SWEEP (John et al., 2019), and StarShip (Basu Roy and Tiwari, 2024). In addition, several studies have investigated scheduling workflows in a multi-cloud environment, utilizing serverless services from various providers, such as AWS Lambda, Google Cloud Functions, and IBM Cloud Functions. Examples of such studies include xAFCL (Ristov et al., 2023), XFaaS (Khochare et al., 2023), SwitchFlow (Chen et al., 2024a), and FaaSt (Ristov and Gritsch, 2022).
However, there are some inherent characteristics of serverless computing that make running workflows difficult or inefficient (Roy et al., 2022). One key feature is the separation of computation and storage, which is designed to support auto-scaling for fluctuating workloads. As a result, serverless platforms use stateless functions that do not retain temporary state information in persistent storage, which creates challenges for workflows that require data storage and exchange between tasks. To address this, brokers must separately provision storage services such as AWS S3 or EFS for data exchange, but these options are slower and less efficient than direct data transfers between tasks (Li et al., 2023a). Additionally, serverless platforms are optimized for lightweight, short-running tasks, and typically impose task execution limits (e.g., 300 seconds for AWS Lambda and 60 seconds for Google Cloud Functions). This makes it difficult to run long-running, high-performance scientific workflows. Furthermore, serverless platforms use a scale to zero approach, where function containers are de-provisioned after a period of inactivity (e.g., 15 minutes in AWS). The cold start issue arises when a new function instance must be launched to handle an incoming request, resulting in delays. This cold start delay can be particularly problematic for short-running tasks, making execution inefficient.
These challenges have led researchers to consider hybrid execution strategies, combining serverful and serverless platforms to run scientific workflows (Malawski et al., 2020). In these hybrid platforms, the issues faced by serverful platforms, such as underutilization of leased VMs, are mitigated by the use of serverless platforms, and vice versa. In hybrid approaches, workflows are executed on a combination of traditional serverful (using VMs) and modern serverless (using functions) cloud platforms. VMs are well-suited for long-running or sequential tasks, while functions are more appropriate for lightweight, parallel tasks. Additionally, the local storage of the VM can be leveraged to share large datasets between tasks. When there is a significant volume of data to be transferred between two tasks, scheduling them on the same VM can eliminate data transfer delays, as they can use shared storage. Serverless services can be paired with any serverful pricing model (on-demand, spot, or reserved) to reduce execution costs and makespan. Although a few researchers have considered scheduling scientific workflows on hybrid platforms, most of them focus on scheduling a single workflow.
Roy et al. (Roy et al., 2022) proposed Mashup, a framework for executing scientific workflows in a hybrid environment that combines VMs and functions. They formulated an optimization problem to determine the optimal execution environment for each task in a workflow, aiming to minimize the overall execution time. The solution involved executing each task on its preferred platform (either a function or a VM). Zhang et al. (Zhang et al., 2024) presented the Budget-Constrained Workflow Scheduling (BCWS) algorithm, a heuristic approach for scheduling a single workflow with a limited budget on a hybrid setup of on-demand VMs and serverless functions, aiming to minimize makespan. The algorithm follows three phases. In the first phase, Initial Schedule Construction, the cheapest functions are allocated to all tasks, and an initial critical path is computed. If any budget remains, faster functions are assigned to critical path tasks to reduce the makespan, continuing until the budget is exhausted. In the second phase, Resource Replacement, the algorithm attempts to replace the rented functions with cheaper but non-slower VMs to restore the remaining budget. The third phase, Budget Reallocation, seeks to assign faster functions using a similar approach to the first phase, further decreasing the makespan.
Other platforms such as Lithops (Eizaguirre et al., 2024), DEWE v3 (Jiang et al., 2017), and HyperFlow on AWS Lambda and ECS (Elastic Container Service) (Balis et al., 2020) have also explored executing workflows in a hybrid environment combining serverful and serverless platforms. However, these platforms focus solely on executing a single workflow, which, while promising, is insufficient for developing a WaaS broker. The only prior work that addresses the execution of multiple workflows in a hybrid environment is Sun et al. (Sun et al., 2024), which is included in this SLR. The integration of serverless FaaS platforms with various pricing models (reserved, spot, and on-demand) for serverful VMs or containers presents a highly promising approach for provisioning resources, aiming to reduce costs and enhance performance. As a result, new provisioning and scheduling algorithms are needed to effectively leverage these emerging serverless platforms for scientific workflow execution.
5.4. Monitoring and Fault Tolerance
The Monitoring and Fault Tolerance component has also received limited attention from researchers. A combination of various fault tolerance methods can be applied in different scenarios. For instance, in a broker using reserved instances, the replication method can be employed when there are a high number of idle reserved instances. Alternatively, the re-execution method can be used for short-running tasks, while checkpointing can be applied for long-running tasks. Additionally, this component could work in conjunction with the negotiation and pricing component to offer different options to the user. For example, users could opt for a more reliable service for executing critical workflows by paying a higher fee. In such a scenario, the pricing component could calculate the cost of task replication, with assistance from the fault tolerance component, and incorporate it into the overall pricing.
6. Conclusion

### v2

@@ 4.4 Monitoring and Fault Tolerance Taxonomy
The final method is replication, which involves running duplicate instances of each task across multiple VMs to enhance the reliability of task execution [161]. In the event of a failure, the task can continue running without disruption, as its duplicate instances remain operational on other resources. While this approach ensures task continuity without losing time, it adds additional costs for the broker due to the need for extra resources to run the duplicated tasks. Chen et al. [38] combine replication and re-execution for workflows running on a mix of on-demand and spot instances. They schedule tasks on spot instances only if there is sufficient time for re-execution in case the resource is reclaimed by the provider. Otherwise, the task is scheduled on an on-demand instance. When tasks are assigned to spot instances, they attempt to identify idle time on on-demand instances to replicate the tasks where possible. If a spot instance is reclaimed, tasks running on it are checked for duplicates. Tasks with duplicates are ignored, while those without duplicates are re-executed on another instance.
4.5 Insights from Previous Work: Trends and Trade-offs
This section provides a concise overview of the insights derived from previous studies. Figure 10 illustrates the relationship between workflow arrival patterns and the scheduling algorithms and types reported in the literature. For online arrival patterns, a broad range of algorithms has been applied, with heuristic and RL-based methods being the most prominent. This aligns with the reliance on dynamic scheduling to handle the unpredictable and spontaneous nature of online workflows. In contrast, batch arrivals are predominantly associated with static scheduling, especially the combined type, which leverages prior knowledge of workflow structures and task execution times. Such scenarios also favor meta-heuristic algorithms, as they are well-suited to exploring the vast search space created by multiple workflows.
Figure 10: Number of Articles by Workflow Arrival Patterns vs. Scheduling Algorithms, and Scheduling Types
(a) Scheduling Algorithms vs. Scheduling Types
(b) Scheduling Algorithms per Year
Figure 11: Number of Articles by Scheduling Algorithms vs. Scheduling Types, and Scheduling Algorithms per Year
Figure 11(a) summarizes the distribution of scheduling algorithms across different scheduling types. As shown, heuristic algorithms have been used in both static and dynamic scheduling, with a notable emphasis on dynamic-immediate strategies. In contrast, meta-heuristic algorithms appear exclusively in static scheduling, while RL-based techniques are predominantly applied in dynamic solutions. Figure 11(b) further illustrates the temporal trend in algorithm usage, revealing a growing interest in RL, alongside meta-heuristics, particularly in recent years—even though the dataset extends only to the first half of 2025.
To conclude, since scheduling algorithms form the core of WaaS brokers and directly impact their overall performance, we provide a comparative evaluation in Table 1. This table highlights the trade-offs among different scheduling algorithms, including their strengths and weaknesses, their suitability for static or dynamic contexts, and considerations such as execution complexity and scalability.
Algorithm Advantages Disadvantages Static/Dynamic Complexity
Heuristic Fast and lightweight,
Suitable for real-time environments,
Scales well to large workflows.
Typically produces suboptimal schedules,
Limited adaptability to dynamic changes,
Lacks generalizability.
Suitable for both Very low complexity,
Typically linear or near-linear
Meta-heuristic Good global exploration,
Effective for complex, multi-objective problems
Capable of exploring large spaces.
High computational cost,
Slow convergence,
High parameter sensitivity.
Mostly static Moderate to high complexity,
Scales poorly with tasks/VMs
Hyper-heuristic Learns from past runs,
Promotes generality and reusability,
Combine strengths of multiple heuristics.
Long training time,
Complex design,
May overfit to training scenarios.
Static training, dynamic scheduling Very high during training,
Scheduling is very fast
Reinforcement Learning Learns from environment interactions,
Adapts well to dynamic and uncertain conditions,
Learns long-term trade-offs.
Long training time,
Reward and feature engineering needed,
Static training, usually dynamic scheduling High training cost, but scheduling is fast and scalable
MCDM Transparent and interpretable,
Fast runtime,
Supports explicit trade-off modeling.
Limited scalability to large decision spaces,
Weight assignment can be subjective,
Typically used only for VM selection.
Usually dynamic. Low complexity.
Game Theory Models interactions among competing objectives,
Offers equilibrium-based fairness and stability,
Models decentralized or competitive scenarios.
High mathematical complexity,
Requires accurate utility modeling.
Usually static High complexity,
Not scalable beyond few players.
Mathematical Optimization Theoretical optimality,
Transparent formulation,
Can encode deadlines and resource constraints precisely.
Difficult to scale to large instances,
Solver overhead,
May not be real-time feasible.
Usually static Very high complexity
Table 1: Comparative evaluation of scheduling algorithms.
5 Future Directions
Despite significant advancements in WaaS brokers, numerous open challenges and research opportunities remain. The rise of artificial intelligence (AI) and large language models (LLMs) opens new possibilities for transforming the design and operation of WaaS brokers through intelligent decision-making, predictive resource management, and automated negotiation, potentially rendering many traditional methods obsolete. Emerging technologies, including serverless and quantum computing, further expand opportunities to enhance broker performance and flexibility. Additionally, underexplored areas such as dynamic negotiation strategies and fault-tolerant mechanisms provide promising directions for improving user satisfaction and overall system robustness. This section highlights these opportunities, aiming to guide researchers toward the most impactful avenues for advancing WaaS broker design and implementation.
5.1 AI-Driven WaaS Brokers
Recent advances in AI, particularly in generative LLMs and DRL, are transforming the landscape of cloud computing. WaaS brokers enhanced with these technologies could harness them throughout the workflow lifecycle, potentially redefining both their design and performance.
DRL enables dynamic decision-making in uncertain and continuously changing cloud environments. By learning optimal policies for task scheduling and resource provisioning through direct interaction with the environment, DRL-based brokers can adapt to workload fluctuations, heterogeneous resources, and shifting performance goals. This adaptability makes scheduling decisions more efficient and resilient. Figure 11(b) illustrates that reinforcement learning has already received strong research attention for workflow scheduling in recent years, and its potential for further applications remains significant.
Meta-RL [23] extends these capabilities by enabling brokers to quickly adapt to new workflow scenarios with minimal experience. Instead of tailoring policies only for specific environments, meta-RL learns a meta-policy that generalizes across many scenarios. In the context of multiple workflows, this means the scheduler can leverage knowledge gained from previously encountered workflow structures, arrival patterns, and resource conditions to make faster and more effective decisions for unseen workloads. During training, brokers are exposed to synthetic but diverse situations—such as varying workflow DAGs, fluctuating resource availability, or dynamic pricing. As a result, the meta-policy learns general adaptation strategies rather than memorizing solutions for specific cases, making it especially suitable for multi-tenant WaaS environments where workflow types, deadlines, and system conditions can change unpredictably.
LLMs bring a different but complementary strength. They introduce adaptive, context-aware reasoning into scheduling and provisioning decisions, which were traditionally driven by fixed rules or heuristics. An LLM-enhanced broker can parse high-level workflow descriptions, SLAs, and performance objectives expressed in natural language, translating them into executable scheduling and provisioning plans. It can analyze task dependencies, predict execution times from logs, and choose suitable resource types, whether serverless or traditional VMs, and pricing models. With access to real-time monitoring data, LLMs can also detect SLA violations or anomalies, identify root causes, and recommend or directly trigger corrective actions such as scaling resources or rescheduling tasks. Beyond automation, LLMs can serve as a natural interface between users and brokers, allowing administrators to specify goals in plain language—for example: ”allocate spot instances to flexible tasks”—which the broker then translates into scheduling directives.
Research in this area is still nascent. While no studies currently apply LLMs directly to workflow scheduling, a few early works explore their use in related contexts. Jadhav et al. [79] introduced ReAct (Reason+Action), an LLM-based scheduler for multi-objective job scheduling in HPC environments. Wu et al. [183] proposed a hierarchical language-agent framework that generates context-aware VM scheduling policies using LLM reasoning capabilities to schedule VMs on physical machines. Similarly, Akbari et al. [6] introduced IntentContinuum, an LLM-based framework for intent-driven resource management across edge–cloud systems. User goals (e.g., minimizing latency or energy use) are expressed as intents, which the framework monitors against system performance. When violations occur, the LLM identifies root causes and suggests corrective actions through structured templates.
In many cases, the general reasoning capabilities of state-of-the-art LLMs are sufficient to generate effective scheduling decisions when these inputs are clearly provided in natural language or structured formats. However, for scientific workflows—especially multiple concurrent workflows—scaling such approaches requires grounding LLMs with domain knowledge. Retrieval-Augmented Generation (RAG) can improve accuracy by incorporating historical workflow traces or performance data, while fine-tuning can help internalize specialized scheduling strategies. Together, these techniques improve efficiency, consistency, and adaptability, enabling brokers to balance cost, performance, and deadline requirements in highly dynamic environments.
Another promising direction is the integration of LLMs and RL for scheduling and provisioning in WaaS brokers. In this setup, the LLM functions as a high-level strategist, generating candidate scheduling and provisioning plans by reasoning over workflow deadlines, dependencies, available resources, and pricing schemes. These plans are then passed to an RL agent, which interacts with the environment to evaluate their effectiveness against performance metrics such as makespan, cost, energy efficiency, and SLA compliance. Through continuous feedback, the RL agent refines and adapts the strategies, ensuring they are well-tuned to real-world dynamics. This synergy combines the broad reasoning and contextual awareness of LLMs with the optimization power of RL, enabling brokers to be both innovative in planning and robust in execution. While this approach has not yet been applied to workflow scheduling specifically, similar LLM–RL frameworks have shown promise for independent task scheduling [164, 93].
5.2 Toward Intelligent Negotiation and Autonomous Fault Tolerance
Two dimensions that remain surprisingly underexplored in WaaS broker research are negotiation/pricing mechanisms and fault tolerance strategies. Yet, these components hold the potential to redefine the role of the broker from reactive middleware into autonomous, market-aware, and self-healing ecosystems.
On the negotiation side, most existing approaches still depend on static pricing models and rigid SLAs, which do not capture the dynamic nature of multi-tenant cloud environments. Future brokers could instead employ predictive, multi-objective negotiation engines that continuously adapt SLAs to workload demands, resource availability, and market conditions. By computing Pareto-optimal solutions in multi-criteria optimization settings, brokers could present users with structured QoS–price trade-off options (e.g., higher reliability at a higher cost). Furthermore, pricing strategies should evolve from simple cost-based formulas to value-driven, market-oriented models [182], incentivizing behaviors such as off-peak submissions when reserved resources are underutilized or flexible deadlines that take advantage of spot instances. In this vision, the broker acts as a market strategist, harmonizing user and provider incentives while fostering both efficiency and fairness.
LLMs reinforce this vision by enabling adaptive, human-like negotiation. Unlike traditional rule-based, LLMs can engage in conversational bargaining with users, interpreting natural concerns such as ”the cost is too high”, and proposing trade-offs such as offering a reduced price in exchange for a longer completion time. LLMs are also able to analyze market conditions, compare pricing models, and take SLA constraints into account when predicting the likely cost of execution and generating adaptive offers or counteroffers. By retrieving historical negotiation outcomes and pricing data via RAG, brokers can further ground their strategies in real-world patterns, making the negotiation process more efficient, transparent, and user-friendly. Recent work on negotiation between humans and software agents using LLMs [25, 142, 168] highlights promising directions that could be directly adopted in WaaS broker negotiation frameworks.
Fault tolerance is another area where a fundamental rethinking is necessary. At present, brokers largely depend on reactive strategies like retries or failovers. Future designs should combine predictive modeling, hybrid strategies, and proactive resilience. Depending on the context, replication could leverage idle reserved instances, re-execution could handle short tasks, and checkpointing could protect long-running jobs. DRL could be leveraged to identify tasks that are critical or highly fault-prone, and to determine whether they should be duplicated or checkpointed to minimize the impact of failures. Crucially, this layer could be intertwined with negotiation and pricing, enabling users to select reliability levels as a priced option—turning fault tolerance itself into a negotiable, market-integrated feature. More radically, by combining digital twins [78] of datacenters with survival analysis and deep learning, brokers could simulate cascading failure scenarios in real time, migrating or replicating workflow components before bottlenecks or SLA violations occur. Over time, such systems would refine their resilience through continual feedback loops, effectively learning what-if strategies for complex multi-tenant environments.
LLMs can also significantly enhance monitoring and fault-tolerance functions by converting raw telemetry into actionable insights. Unlike traditional monitoring systems that rely on rule-based anomaly detection, LLMs can analyze heterogeneous inputs such as logs, metrics, error codes, and workload traces to identify anomalies, determine the root cause of performance degradations, and provide human-readable diagnoses. In the context of fault tolerance, LLMs can reason about appropriate recovery strategies—such as retrying transient failures, migrating tasks to alternative resources, duplicating critical tasks, or triggering checkpoints—based on workflow topology, task criticality, and SLA constraints. Moreover, LLMs can learn or retrieve patterns from historical execution data to guide proactive decisions; for example, ”for DAGs with long critical paths, prioritize checkpointing of bottleneck tasks.” By integrating such knowledge through RAG, LLMs can ground their reasoning in domain-specific logs, past workflow traces, and known fault scenarios, ensuring more accurate and context-aware recommendations, such as avoiding unreliable spot instances for deadline-sensitive workflows.
Traditionally, DevOps engineers handled fault detection, localization, and mitigation. The concept of AI for IT Operations (AIOps) has been proposed as a means of automating these processes by applying machine learning to the large volumes of monitoring and infrastructure data, thereby enabling automatic detection, diagnosis, and resolution of system issues with minimal human involvement [152]. Recent advances in AIOps have introduced LLM-based agents that exploit natural language processing to interpret diverse input data and logs, identify causes of faults, and even generate scripts that invoke external tools for automated remediation [152, 189]. Zhang et al. [193] provide a recent survey of LLM-based AIOps systems. Adapting such techniques to WaaS brokers could reduce operational overhead while improving robustness and resilience.
5.3 WaaS Brokers as Cloud Economists
Although some studies have explored combining pricing models (on-demand, reserved, and spot) for WaaS brokers, more sophisticated approaches are still needed. A promising direction is to reconceptualize brokers not just as technical mediators between users and CSPs, but as intelligent cloud economists. By leveraging multiple pricing models, brokers could act as market agents that strategically navigate uncertainty and volatility. Instead of treating pricing models as static choices, future brokers could dynamically orchestrate them based on long-term workload predictions and short-term market signals. For example, they might proactively acquire reserved capacity for recurring workflow patterns, exploit spot markets for opportunistic savings, and rely on on-demand instances to meet deadlines. In this way, brokers function as self-optimizing portfolios of cloud resources, balancing risk, cost, and performance much like financial traders manage assets in dynamic markets.
This vision can be realized through integrating LLMs and RL into a unified decision-making loop. LLMs, with their ability to interpret workflow descriptions, pricing policies, and historical traces, can serve as strategists that map user requirements to high-level provisioning strategies. RL agents, in turn, act as executors, refining these strategies under uncertainty by adapting to workload variability and market fluctuations. With RAG grounding decisions in past negotiation outcomes and cost histories, brokers could evolve into proactive, self-learning agents that bargain with users, forecast demand, and arbitrage across pricing models. WaaS brokers, therefore, evolve beyond simple workflow managers, emerging as autonomous, market-aware entities—cloud economists that redefine how scientific workflows are priced, provisioned, and executed at scale.
5.4 Federated Brokers for Cross-Cloud and Data-Sensitive Workflows
As workflows increasingly span multiple administrative domains or involve sensitive data in domains such as healthcare, finance, and genomics, WaaS brokers must ensure privacy, data locality, and regulatory compliance. Traditional centralized brokers collect the full workflow DAG and raw data at a single control point, which simplifies orchestration but exposes sensitive inputs and may violate legal or organizational constraints. Federated brokers, in contrast, adopt a decentralized architecture where each participating domain or CSP retains control of its data and tasks, contributing only high-level metadata, encrypted updates, or aggregated statistics to the global decision process. Each cloud region or administrative domain hosts a local broker agent, and these agents coordinate—optionally guided by a central WaaS broker—to orchestrate workflows while enforcing local policies.
In such a setting, federated brokers must perform task graph partitioning under policy-based placement constraints. Tasks involving sensitive data can remain on-premise or within a specific trusted cloud, while other tasks can be flexibly scheduled to public clouds for efficiency or cost reduction. To coordinate these distributed decisions, federated brokers can employ federated learning to train predictive models for task runtimes across domains, secure multi-party computation or homomorphic encryption for joint optimization of provisioning, and differential privacy to protect workflow-level analytics. This allows brokers to orchestrate end-to-end workflows across heterogeneous clouds while guaranteeing privacy, compliance, and performance scalability.
5.5 Hybrid Serverless/Serverful Provisioning
@@
The rise of serverless computing and the FaaS model introduces new opportunities for WaaS brokers [180]. In this model, users and brokers no longer manage resources or scheduling directly; instead, the CSP handles provisioning, scaling, and function orchestration. FaaS also offers fine-grained, utilization-based pricing, charging only for actual consumption and thus eliminating idle-resource overheads. While several studies have explored applying FaaS to scientific workflows [113], its adoption remains in the early stages. Existing work mostly targets single workflows, with frameworks such as StepConf [181], Wukong [26], SWEEP [84], and StarShip [21]. Other research has investigated multi-cloud scheduling across multiple CSPs, including xAFCL [134], XFaaS [89], SwitchFlow [32], and FaaSt [133].
Despite these advances, several limitations hinder workflow execution on pure serverless platforms [137]. First, the separation of computation and storage, intended to enable auto-scaling, results in stateless functions that cannot retain intermediate states. This creates inefficiencies for workflows requiring data exchange between tasks, forcing reliance on external storage, which is slower than direct task-to-task transfers [107]. Second, serverless environments are optimized for lightweight, short-running tasks, with execution time limits (e.g., 300s in AWS Lambda, 60s in Google Cloud Functions), making them unsuitable for long-running long-running tasks. Finally, the scale-to-zero mechanism de-provisions idle function containers after a period of inactivity, leading to cold start delays that particularly affect short tasks.
To overcome these issues, hybrid provisioning strategies have been proposed, combining serverful and serverless platforms [113]. In such environments, the weaknesses of each model are offset by the strengths of the other: VMs are suited for long-running or sequential tasks, while functions excel at lightweight, parallel workloads. Local VM storage can facilitate large data exchanges between tasks by eliminating transfer overheads, while serverless functions reduce underutilization and improve elasticity. Serverless services can also be paired with various VM pricing models to lower both cost and makespan. DRL and LLM-based decision engines may further optimize task placement by deciding whether a task should run as a function or on a VM based on execution time, data size, or deadlines, and by selecting the most cost-effective pricing model.
Several studies have investigated hybrid execution, though most focus only on single workflows. Notable examples include Mashup [137], Lithops [50], DEWE v3 [83], HyperFlow on AWS Lambda [19], and the Budget-Constrained Workflow Scheduling (BCWS) algorithm [192]. The only existing work that considers multiple workflows in a hybrid environment is by Sun et al. [157], which is covered in this SMS. Overall, the integration of serverless FaaS with serverful VMs under diverse pricing models offers a highly promising direction for WaaS brokers. Achieving this vision calls for new provisioning and scheduling algorithms capable of fully exploiting hybrid platforms to minimize costs and improve performance in scientific workflow execution.
5.6 Hybrid Quantum/Classical WaaS Brokers
As discussed earlier, quantum computing remains an emerging, rapidly evolving technology, with current systems offering only limited capabilities. Consequently, hybrid HPC/quantum environments—and the corresponding hybrid WMSs—are still in the research and development phase. The architectural and behavioral differences between quantum and classical computers make existing WMSs and their provisioning and scheduling algorithms inadequate for hybrid scientific workflows. Cranganore et al. [44] highlight the main challenges of hybrid workflow execution and propose a high-level architectural design for a hybrid WMS built on Pegasus, detailing essential software components, their roles, and the system-level requirements.
A first major challenge is the heterogeneity of quantum hardware technologies—such as superconducting qubits, trapped ions, and optical/photonics—which differ in number of qubits, supported quantum gates, and coherence times [24]. These variations may render certain algorithms inefficient or even infeasible on specific hardware. Moreover, no standard set of performance descriptors yet exists. While metrics like IBM’s Quantum Volume (QV) and Circuit-Layer Operations per Second (CLOPS) have been introduced, they remain evolving and non-standardized. To cope with this heterogeneity, a hybrid WaaS broker must provide interfaces to multiple quantum backends and support transpilation of high-level circuits into hardware-specific representations—assuming execution is feasible. This requires a flexible programming interface that abstracts hardware details, ensuring portability and adaptability across quantum platforms [149].
A second challenge is developing accurate performance models for predicting execution time, error rates, and other key metrics on specific quantum hardware. Such models are critical for broker schedulers to make informed allocation decisions, yet constructing generalizable models across quantum platforms remains unresolved [44]. Other challenges include hybrid workflow optimization—through techniques like gate-level minimization to reduce circuit depth—and error mitigation in noisy intermediate-scale quantum (NISQ) devices, where noise and decoherence still hampers reliability.
Despite these challenges, some researchers have proposed initial frameworks designed to define and execute single hybrid workflows in quantum/classical environments. Notably, Weder et al. introduced MODULO [179, 178], a hybrid WMS with components for workflow generation, optimization, packaging, deployment, and monitoring. While their evaluation targets a single workflow, the design could extend to multiple workflows. Other efforts include Pilot-Quantum [115], XFaaSQ—a FaaS-based hybrid workflow engine [80], and QCCP (Classical-Quantum Collaborative Computing) [49].
The only current WMS explicitly supporting multiple hybrid workflows is Qonductor [60], still a preprint in mid-2025. Its modular architecture consists of a client, a leader, and one or more workers. The leader node houses the resource estimator, hybrid scheduler, and job manager. The estimator generates alternative resource plans based on workflow fidelity and estimated execution time; clients select one plan, which is then executed by the job manager. The job manager coordinates the execution of ready jobs by invoking the hybrid scheduler to allocate appropriate resources. The hybrid scheduler applies a simple filtering-scoring heuristic for classical jobs, and a three-stage process for quantum jobs—preprocessing, optimization, and selection. During optimization, NSGA-II produces a Pareto front balancing completion time and fidelity, and then an MCDM method selects the best option. Qonductor also includes monitoring and fault-tolerance mechanisms to improve reliability. Importantly, its evaluation demonstrates multiple workflow scheduling and execution—an advance over existing work.
In parallel, researchers at Oak Ridge National Laboratory (ORNL) have developed a robust framework for embedding QC into classical HPC environments [24, 149]. It features a unified resource manager orchestrating both quantum and classical resources, along with a versatile programming interface that abstracts hardware differences. Their proposed architecture is notably rich and addresses a wide range of requirements for a cohesive hybrid environment. It introduces key abstractions for task decomposition and coordination, enabling efficient interaction between quantum and classical tasks. Such designs can be adopted by WMSs like Pilot-Quantum to facilitate hybrid workflows.
Looking ahead, research should focus on developing hybrid quantum/classical WaaS brokers with provisioning and scheduling algorithms tailored for quantum workloads. These brokers must consider quantum-specific challenges, including diverse performance descriptors and hardware-dependent performance models, to achieve efficient, scalable hybrid workflow execution.
6 Conclusion